# AI-Based Vulnerability Analysis of NFT Smart Contracts

### v2

@@ 1. INTRODUCTION
NFT (Non-fungible Token), or non-fungible token, is a certified storage unit operating on a unique and indivisible blockchain platform. Currently, it is mostly in the form of electronic files of the source files of artistic creation, and its value is reflected by virtual cryptocurrencies(Kong et al., 2024).
@@
The current focus on the NFT smart contract market revolves mainly around four aspects: classification and mining of vulnerabilities, development of audit tools, repair of vulnerability strategy, and construction of the NFT ecosystem(Liu and Li, 2025). Due to the complexity of NFT smart contract vulnerabilities, there is currently no mature large-scale artificial intelligence detection model specifically for NFT smart contract vulnerabilities in the market.(Li et al., 2024b) Therefore, the analysis, detection, repair, and maintenance of vulnerabilities in the NFT smart contract require a significant amount of resources, which is not conducive to the continuous healthy development of the existing NFT market.
@@
Therefore, this paper proposes an AI-based NFT smart contract vulnerability analysis project in response to the lack of large-scale artificial intelligence analysis models. The goal is to analyze vulnerabilities and carefully analyze the generated models, starting from data, models, and effects, and conducting a large amount of demonstration and experimental work (Li et al., 2024f). We strive to accumulate experience in NFT smart contract vulnerabilities and contribute to the application of large-scale artificial intelligence models.
@@
There are already several examples of defect detection in AI-based smart contracts,rocessing (NLP) and machine learning algorithms to perform static analysis and vulnerability detection on smart contract code(Bu et al., 2025), identifying potential security risks and providing improvement suggestions, offering reliable security assurance for blockchain developers;(Niu et al., 2024) Li Tao and others proposed a public audit of smart contracts based on game theory; Chuang Ma1 proposed HGAT,(Ma et al., 2023) a hierarchical graph attention network-based detection model. Internationally, the Harvard University Blockchain Security Laboratory has developed a static analysis tool for smart contracts using artificial intelligence technology, which can automatically detect vulnerabilities and security risks in smart contracts; and Anzhelika Mezina (Mezina and Ometov, 2023) and others proposed a method combining binary classification and multiclassification to detect vulnerability in smart contracts in their paper.
@@
This paper focuses on the currently widely used NFT smart contracts, conducting a comprehensive and in-depth study on their related security issues, collecting a large number of applied NFT smart contract codes; after careful sorting and differentiation, the attacks are divided into five categories; they are processed and studied separately and finally trained, in addition, this paper also continues to explore its shortcomings and possible future work directions and priorities, providing its own opinions for the improvement of NFT smart contract detection technology.
@@
  • Sufficient and accurate data: This paper compiled a dataset of 16,527 smart contracts, and analyzed and labeled the vulnerabilities in their line codes. We also carefully cleaned and preprocessed the data to exclude inaccurate or low-quality code samples, ensuring that the data quality on which the model is based is highly reliable.
  • Targeted solutions: We conducted an in-depth analysis of NFT smart contracts and discovered five main defects: Risky Mutable Proxy, ERC-721 Reentrancy, Unlimited Minting, Missing Requirements, and Public Burn. These defects may seriously impact the contract’s security and robustness. To solve these problems, we formulate targeted solutions to improve the program’s quality and robustness.
  • Sufficient training and organization: In our research, we performed a detailed parameter tuning for the random forest model. Adjusting the number of trees, the depth of trees, the selection of features, and other parameters, we explored many different training results. This sufficient training and organization process improved our model and provided more reliable prediction and evaluation capabilities for smart contract code security.
@@ 2.2. Random Forest Model
In random forests, when we need to predict new datasets, to ensure accuracy, each decision tree independently predicts the sample without interference from other decision trees. For classification problems, the final prediction result can be determined through a one-vote-per-tree voting process, with the majority vote determining the best result. For regression problems, we sum all predicted values and average them to eliminate errors, ultimately obtaining the result.
In summary, the random forest model is based on the construction of multiple decision trees using random data sampling and random feature selection methods. Through continuous iteration, it integrates the prediction results of each decision tree, thereby improving the accuracy and adaptability of the model.
@@ 4. CONSTRUCTION OF DECISION TREES
We will address the five vulnerabilities listed above, perform feature extraction, handle training, and finally, summarize the analysis and application for general models.
@@ 4.1. Feature Extraction
  • Risk Mutable Proxy Feature Extraction:For the risk of Mutable proxy vulnerability, feature extraction is conducted from six aspects: whether there is a proxy contract call, whether there is a parameter setting function, whether there is permission control, whether there is a state change record, whether there is a risk operation function, and whether there is condition detection.
  • ERC-721 Reentrancy Feature Extraction:For the ERC-721 reentrancy vulnerability, feature extraction is performed from eight aspects: whether there is an external call, whether there is a locking mechanism in the contract, whether the contract state is not properly handled, whether there is an asset transfer vulnerability in the contract, whether the contract state modification is separated, whether the contract does not properly handle the return value of external contract calls, whether the contract lacks necessary state checks and updates, and whether the contract does not properly handle exceptions.
  • Unlimited Mining Feature Extraction:For unlimited mining vulnerability, feature extraction is conducted from eight aspects: whether the minting function lacks permission verification, whether there is a lack of total supply limit, whether there is a lack of condition checks, whether there is an extensible minting logic, whether there is a lack of a triggerable stop mechanism in the smart contract, whether there is an unreasonable minting fee in the smart contract, whether there are external dependency calls in the smart contract code, and whether there is a possibility of unlimited minting due to permission or role abuse in the contract.
  • Missing Requirement Feature Extraction:For the Missing Requirement vulnerability, feature extraction is conducted from eight aspects: whether there is a function definition in the contract but lacks an input validation function, whether there is a lack of security checks, whether there is a lack of transfer restriction functions, whether there is a lack of auditing and readability functions, whether there is a lack of event triggering, whether there is a lack of permission control, whether there is an upgrade mechanism, and whether there is a lack of asset metadata validation.
  • Public Burn Feature Extraction:For the public Burn vulnerability, feature extraction is conducted from six aspects: whether there is a lack of authentication, whether there is a lack of confirmation or recovery mechanism, whether there is a mismatched authorization mechanism, whether the owner identity is not considered, whether there is a lack of event logs, and whether there is a duplicate destruction mechanism.
@@ 4.2. Feature Selection and Division
  • The ID3 algorithm lacks a pruning step for the generated decision tree, which can easily lead to overfitting (Lalduhsaka et al., 2022).
  • The C4.5 algorithm requires sorting numerical category attributes during tree construction, which requires ensuring that the generated data can be stored in the host memory’s dataset [18]. When the provided data volume is too large, the generated data volume becomes too large, making it difficult to run the program.
@@
  • The CART algorithm uses the Gini criterion for judgment during training data processing, as it does not require logarithmic operations that consume a lot of resources (Chen et al., 2023). Given the large data volume in this study, it is more suitable.
@@ 4.5. Decision Tree Pruning
  • When $\alpha$=+$\infty$+ ∞, the regularization degree is very high, meaning the decision tree containing only the root node of the CART tree is the best-performing subtree. Generally, the larger $\alpha$ is, the more thorough the pruning, and the better the effect.
  • Using a recursive method, starting from zero, $\alpha$ increases sequentially, $\alpha_{0}$¡$\alpha_{1}$$\ldots$$\alpha_{n-1}$¡$\alpha_{n}$¡+$\infty$+ ∞,
@@
When $\alpha=0$ or $\alpha$to $+0$
@@
In this paper, g(t) represents the degree of reduction in the overall loss function after pruning. For example: in $T_{0}$, pruning the $T_{t}$ with the smallest g(t) value, the resulting subtree is $T_{1}$, and this g(t) value is set as $\alpha_{1}$. We get $T_{1}$ as the optimal subtree for the interval [$\alpha_{1}$,$\alpha_{2}$).
Then, iterate until the root node, forming a sequence of { $T_{0}$,$T_{1}$$\ldots$,$T_{n}$. Using the Gini index criterion mentioned in 3.3, test the subtree sequence on the new validation set, select the best-performing subtree, and output it as the optimal decision tree.
@@ 5.1. Introduction to Ensemble Learning
The Random Forest model, as an ensemble learning algorithm, is based on weak classifiers. When dealing with classification and regression problems, the final results are processed through voting and averaging methods (Pawar et al., 2023), ensuring the accuracy and adaptability of the overall model. Due to its excellent stability, it is widely used in various business scenarios.
@@
Ensemble learning is not a specific step or algorithm but rather a concept. We can use the stories ”Many hands make light work” and ”Three Cobblers with their wits combined surpass Zhuge Liang” as examples. Ensemble learning leverages the principle of ”many hands make light work.” It does not create something new like the cobblers but integrates existing algorithms to improve accuracy (Yu et al., 2010). In terms of completing tasks, the approach can be divided into three categories: Stacking, Boosting, and Bootstrap Aggregating (Bagging).
@@
  • Bagging: Bagging (Bootstrap Aggregating) uses the bootstrap method to draw a large number of samples from the original dataset for training. After training, the samples are returned, and multiple independent base learners are trained using these subsamples. Finally, their prediction results are processed to generate the final ensemble model. The core idea of bagging is voting (Rout et al., 2018). Each model has equal influence, and the final answer is determined by voting. Typically, the results obtained through bagging have a smaller variance.
@@
    The main idea of boosting is to appropriately weight the performance of the data in each round of operation. In each iteration, the learner’s weights are readjusted. The data is processed again with incomplete resampling, allowing the data to be retrained, and then their weights are updated based on their performance (Nakach et al., 2022). Through this method, the model reduces the influence of problematic learners, thereby improving overall performance.The general process of boosting is as follows:
@@ 5.1.2   Core Issues in Ensemble Learning
    Stacking: First, multiple individual learners (of different types) are stacked together, then trained, and combined with a meta-learner to predict results. This method allows for better utilization of diverse learners.
@@ 5.3.1   Advantages
Finally, Random Forest shows strong robustness to missing features, maintaining high prediction accuracy. In summary, the Random Forest model is a powerful and comprehensive machine learning algorithm suitable for handling high-dimensional, dense data.
@@ 5.3.2   Disadvantages
Therefore, better handling of abnormal problems and improving the model’s interpretability require further consideration and processing. (Usually, careful discrimination is needed in data processing.) Additionally, for problematic decision trees, providing reasonable and effective explanations, analyzing the causes of problems, and verifying and correcting them are necessary to improve the model’s interpretability.
@@ 5.4. Parameter Tuning
Our Random Forest model has five main parameters: maximum number of features, number of trees, maximum depth, minimum samples required to split an internal node, and minimum samples required at a leaf node. Initially, we do not know where to apply these parameters to achieve optimal results, which requires extensive processing and experimentation. This is where parameter tuning comes into play. Through parameter tuning, we can improve the model’s robustness and accuracy in handling complex problems, ensuring its normal operation on general issues.
@@ 5.4.2   GridSearch Hyperparameter Tuning Demonstration
GridSearch is a hyperparameter tuning algorithm we learned in our artificial intelligence and big data course. It allows us to control a model’s complexity through hyperparameters, thereby influencing aspects such as the model’s accuracy and training time.

### v3

@@ 1. INTRODUCTION
NFT (Non-fungible Token), or non-fungible token, is a certified storage unit operating on a unique and indivisible blockchain platform. Currently, it is mostly in the form of electronic files of the source files of artistic creation, and its value is reflected by virtual cryptocurrencies(Kong et al., 2024; Huang et al., 2023).
@@
The current focus on the NFT smart contract market revolves mainly around four aspects: classification and mining of vulnerabilities, development of audit tools, repair of vulnerability strategy, and construction of the NFT ecosystem(Liu and Li, 2025). Due to the complexity of NFT smart contract vulnerabilities, there is currently no mature large-scale artificial intelligence detection model specifically for NFT smart contract vulnerabilities in the market.(Li et al., 2024b) Therefore, the analysis, detection, repair, and maintenance of vulnerabilities in the NFT smart contract require a significant amount of resources, which is not conducive to the continuous healthy development of the existing NFT market. (Gandhi, 2025)
@@
Therefore, this paper proposes an AI-based NFT smart contract vulnerability analysis project in response to the lack of large-scale artificial intelligence analysis models. The goal is to analyze vulnerabilities and carefully analyze the generated models, starting from data, models, and effects, and conducting a large amount of demonstration and experimental work (Li et al., 2024f). We strive to accumulate experience in NFT smart contract vulnerabilities and contribute to the application of large-scale artificial intelligence models. (Praitheeshan et al., 2019)
@@
There are already several examples of defect detection in AI-based smart contracts,rocessing (NLP) and machine learning algorithms to perform static analysis and vulnerability detection on smart contract code(Bu et al., 2025), identifying potential security risks and providing improvement suggestions, offering reliable security assurance for blockchain developers;(Niu et al., 2024) Li Tao and others proposed a public audit of smart contracts based on game theory; Chuang Ma1 proposed HGAT,(Ma et al., 2023) a hierarchical graph attention network-based detection model. Internationally, the Harvard University Blockchain Security Laboratory has developed a static analysis tool for smart contracts using artificial intelligence technology, which can automatically detect vulnerabilities and security risks in smart contracts; (Sendner et al., 2023) and Anzhelika Mezina (Mezina and Ometov, 2023) and others proposed a method combining binary classification and multiclassification to detect vulnerability in smart contracts in their paper.
@@
This paper focuses on the currently widely used NFT smart contracts, conducting a comprehensive and in-depth study on their related security issues, collecting a large number of applied NFT smart contract codes; after careful sorting and differentiation, the attacks are divided into five categories; they are processed and studied separately and finally trained, (Elghalhoud et al., 2022) in addition, this paper also continues to explore its shortcomings and possible future work directions and priorities (Šutas et al., 2024), providing its own opinions for the improvement of NFT smart contract detection technology.
@@
  • Sufficient and accurate data: This paper compiled a dataset of 16,527 smart contracts, and analyzed and labeled the vulnerabilities in their line codes. We also carefully cleaned and preprocessed the data to exclude inaccurate or low-quality code samples, ensuring that the data quality on which the model is based is highly reliable. (Ali et al., 2023)
  • Targeted solutions: We conducted an in-depth analysis of NFT smart contracts and discovered five main defects: Risky Mutable Proxy, ERC-721 Reentrancy, Unlimited Minting, Missing Requirements, and Public Burn. These defects may seriously impact the contract’s security and robustness. To solve these problems, we formulate targeted solutions to improve the program’s quality and robustness. (Yang et al., 2023)
  • Sufficient training and organization: In our research, we performed a detailed parameter tuning for the random forest model. Adjusting the number of trees, the depth of trees, the selection of features, and other parameters, we explored many different training results. This sufficient training and organization process improved our model and provided more reliable prediction and evaluation capabilities for smart contract code security. (Chen et al., 2024)
@@ 2.2. Random Forest Model
In random forests, when we need to predict new datasets, to ensure accuracy, each decision tree independently predicts the sample without interference from other decision trees. For classification problems, the final prediction result can be determined through a one-vote-per-tree voting process, with the majority vote determining the best result. For regression problems, we sum all predicted values and average them to eliminate errors, ultimately obtaining the result. (Boruah et al., 2023)
In summary, the random forest model is based on the construction of multiple decision trees using random data sampling and random feature selection methods. Through continuous iteration, it integrates the prediction results of each decision tree, thereby improving the accuracy and adaptability of the model. (Chen et al., 2024)
@@ 4. CONSTRUCTION OF DECISION TREES
We will address the five vulnerabilities listed above, perform feature extraction, handle training, and finally, summarize the analysis and application for general models.(Loh, 2011)
@@ 4.1. Feature Extraction
  • Risk Mutable Proxy Feature Extraction:For the risk of Mutable proxy vulnerability, feature extraction is conducted from six aspects: whether there is a proxy contract call, whether there is a parameter setting function, whether there is permission control, whether there is a state change record, whether there is a risk operation function, and whether there is condition detection. (Utgoff, 1989)
  • ERC-721 Reentrancy Feature Extraction:For the ERC-721 reentrancy vulnerability, feature extraction is performed from eight aspects: whether there is an external call, whether there is a locking mechanism in the contract, whether the contract state is not properly handled, whether there is an asset transfer vulnerability in the contract, whether the contract state modification is separated, whether the contract does not properly handle the return value of external contract calls, whether the contract lacks necessary state checks and updates, and whether the contract does not properly handle exceptions. (Salzberg, 1994)
  • Unlimited Mining Feature Extraction:For unlimited mining vulnerability, feature extraction is conducted from eight aspects: whether the minting function lacks permission verification, whether there is a lack of total supply limit, whether there is a lack of condition checks, whether there is an extensible minting logic, whether there is a lack of a triggerable stop mechanism in the smart contract, whether there is an unreasonable minting fee in the smart contract, whether there are external dependency calls in the smart contract code, and whether there is a possibility of unlimited minting due to permission or role abuse in the contract. (Luu et al., 2016)
  • Missing Requirement Feature Extraction:For the Missing Requirement vulnerability, feature extraction is conducted from eight aspects: whether there is a function definition in the contract but lacks an input validation function, whether there is a lack of security checks, whether there is a lack of transfer restriction functions, whether there is a lack of auditing and readability functions, whether there is a lack of event triggering, whether there is a lack of permission control, whether there is an upgrade mechanism, and whether there is a lack of asset metadata validation. (Tsankov et al., 2018)
  • Public Burn Feature Extraction:For the public Burn vulnerability, feature extraction is conducted from six aspects: whether there is a lack of authentication, whether there is a lack of confirmation or recovery mechanism, whether there is a mismatched authorization mechanism, whether the owner identity is not considered, whether there is a lack of event logs, and whether there is a duplicate destruction mechanism. (Sürücü et al., 2022)
@@ 4.2. Feature Selection and Division
  • The ID3 algorithm lacks a pruning step for the generated decision tree, which can easily lead to overfitting (Lalduhsaka et al., 2022; Ghaleb and Pattabiraman, 2020).
  • The C4.5 algorithm requires sorting numerical category attributes during tree construction, which requires ensuring that the generated data can be stored in the host memory’s dataset. When the provided data volume is too large, the generated data volume becomes too large, making it difficult to run the program.
@@
  • The CART algorithm uses the Gini criterion for judgment during training data processing, as it does not require logarithmic operations that consume a lot of resources (Chen et al., 2023). Given the large data volume in this study, it is more suitable. (Chen et al., 2020)
@@ 4.5. Decision Tree Pruning
  • When $\alpha$$\infty$, the regularization degree is very high, meaning the decision tree containing only the root node of the CART tree is the best-performing subtree. Generally, the larger $\alpha$ is, the more thorough the pruning, and the better the effect.
[…]
@@
When $\alpha=0$ or $\alpha$ $\to$ +0+0
@@
In this paper, g(t) represents the degree of reduction in the overall loss function after pruning. For example: in $T_{0}$, pruning the $T_{t}$ with the smallest g(t) value, the resulting subtree is $T_{1}$, and this g(t) value is set as $\alpha_{1}$. We get $T_{1}$ as the optimal subtree for the interval $\alpha_{1}$$\alpha_{2}$).
[…]
@@ 5.1. Introduction to Ensemble Learning
The Random Forest model, as an ensemble learning algorithm, is based on weak classifiers. When dealing with classification and regression problems, (Sürücü et al., 2022) the final results are processed through voting and averaging methods (Pawar et al., 2023), ensuring the accuracy and adaptability of the overall model. Due to its excellent stability, it is widely used in various business scenarios. (Sagi and Rokach, 2018)
@@
Ensemble learning is not a specific step or algorithm but rather a concept. We can use the stories ”Many hands make light work” and ”Three Cobblers with their wits combined surpass Zhuge Liang” as examples. Ensemble learning leverages the principle of ”many hands make light work.” It does not create something new like the cobblers but integrates existing algorithms to improve accuracy (Yu et al., 2010; Hossain and Islam, 2023). In terms of completing tasks, the approach can be divided into three categories: Stacking, Boosting, and Bootstrap Aggregating (Bagging).
@@
  • Bagging: Bagging (Bootstrap Aggregating) uses the bootstrap method to draw a large number of samples from the original dataset for training. After training, the samples are returned, and multiple independent base learners are trained using these subsamples. Finally, their prediction results are processed to generate the final ensemble model. The core idea of bagging is voting (Rout et al., 2018). Each model has equal influence, and the final answer is determined by voting (Boruah et al., 2023). Typically, the results obtained through bagging have a smaller variance.
@@
    The main idea of boosting is to appropriately weight the performance of the data in each round of operation. In each iteration, the learner’s weights are readjusted. The data is processed again with incomplete resampling, allowing the data to be retrained, and then their weights are updated based on their performance (Nakach et al., 2022). Through this method, the model reduces the influence of problematic learners, thereby improving overall performance. (chiu2022applying) The general process of boosting is as follows:
@@ 5.1.2  Core Issues in Ensemble Learning
    Stacking: First, multiple individual learners (of different types) are stacked together, then trained, and combined with a meta-learner to predict results. This method allows for better utilization of diverse learners. (Togatoropa et al., 2022)
@@ 5.3.1  Advantages
Finally, Random Forest shows strong robustness to missing features, maintaining high prediction accuracy. In summary, the Random Forest model is a powerful and comprehensive machine learning algorithm suitable for handling high-dimensional, dense data. (Dong et al., 2020)
@@ 5.3.2  Disadvantages
Therefore, better handling of abnormal problems and improving the model’s interpretability require further consideration and processing. (Usually, careful discrimination is needed in data processing.) Additionally, for problematic decision trees, providing reasonable and effective explanations, analyzing the causes of problems, and verifying and correcting them are necessary to improve the model’s interpretability. (Li et al., 2022)
@@ 5.4. Parameter Tuning
Our Random Forest model has five main parameters: maximum number of features, number of trees, maximum depth, minimum samples required to split an internal node, and minimum samples required at a leaf node. Initially, we do not know where to apply these parameters to achieve optimal results, which requires extensive processing and experimentation. This is where parameter tuning comes into play. Through parameter tuning, we can improve the model’s robustness and accuracy in handling complex problems, ensuring its normal operation on general issues. (Wang and Xia, 2016)
@@ 5.4.2  GridSearch Hyperparameter Tuning Demonstration
GridSearch is a hyperparameter tuning algorithm we learned in our artificial intelligence and big data course. It allows us to control a model’s complexity through hyperparameters, thereby influencing aspects such as the model’s accuracy and training time. (Yates et al., 2023)